OpenStudy (anonymous):

divide the polynomials (9x 4 + 3x 3y - 5x 2y 2 + xy 3) and (3x 3 + 2x 2y - xy 2)

OpenStudy (anonymous):

\[ \frac{9 x^4+3 x^3 y-5 x^2 y^2+x y}{3 x^3+2 x^2 y-x y^2}=\frac{x (3 x-y)^2 (x+y)}{x (3 x-y) (x+y)} \] Cancel and you are done

OpenStudy (anonymous):

thank you!